OpenStudy (anonymous):

Solve the following system of equations. x - y = 16 2x + y = 2 Help me, please.

OpenStudy (anonymous):

x - y = 16 + y + y x = 16 + y 2(16 + y) + y = 2 32 + 2y + y = 2 32 + 3y = 2 -32 -32 3y = (-30) -- ---- 3 3 y = (-10)

OpenStudy (anonymous):

x - y = 16 x - 10 = 16 + 10 + 10 x = 26

OpenStudy (anonymous):

So, the solution is: (26, -10) Is that one of your answer choices?

OpenStudy (anonymous):

there is no choices, its one of those written responces.

OpenStudy (anonymous):

Wait, don't turn that last answer in. I messed up!!!

OpenStudy (anonymous):

okay.

OpenStudy (amistre64):

....just as a side note tho; x not equal 26

OpenStudy (amistre64):

what methods do you know of to solve a system of equations?

OpenStudy (anonymous):

substitution.. a little bit.

OpenStudy (amistre64):

substitution is fine, but its sooo labor intensive. I perfer elimination

OpenStudy (amistre64):

notice that the y values have different signs, which means if we add these equations together, we cancel out the ys x - y = 16 2x + y = 2 ----------- 3x + 0 = 18 can you solve for x now?

OpenStudy (anonymous):

if i were to do elimination, would i multiply the top equation by 2?

OpenStudy (amistre64):

you could multiply the top by -2, that would be suitable; but the ys are already primed for elimination :)

OpenStudy (anonymous):

x=6?

OpenStudy (amistre64):

correct, x=6

OpenStudy (amistre64):

since x=6 lets use the top, since its structure is simpler x - y = 16 ; x=6 6 - y = 16 can you find y now?

OpenStudy (anonymous):

-y=10. how do you get rid of the -y?

OpenStudy (amistre64):

multiply both sides by -1

OpenStudy (anonymous):

ah, so its y=-10.

OpenStudy (amistre64):

good, now lets check that in the 2nd equation :) 2(6) + (-10) = 2 12 - 10 = 2 ; i like it

OpenStudy (anonymous):

The solution is (6,-10). Sorry for taking so long @alienababy

OpenStudy (amistre64):

John, we perfer to study thru to a solution on this site. Helps people to get more confident in their abilities to solve the problems in general

OpenStudy (amistre64):

but i do agree, that answer is (6,-10) ;)

OpenStudy (anonymous):

so i got the answer ? :)

OpenStudy (anonymous):

Thanks for the hint @amistre64. But my names not John because I'm a girl and I mixed my name and my friends name together. But @alienababy just got stuck and she figured all the other problems out herself.!

OpenStudy (amistre64):

you got it

OpenStudy (anonymous):

Yes, @alienababy. I told you, you were a math genius. You just needed the right motivation!

OpenStudy (anonymous):

thankyou @amistre64 && @Johnjakile1998

OpenStudy (anonymous):

any more questions?

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

nope. :)